OpenStudy (cometailcane):

A liquid helium storage tank has a volume of 0.275 m3; how many liters of liquid helium can the tank hold? Given 1 mL = 1 cm3

OpenStudy (lposton):

Hi

OpenStudy (lposton):

Do u k I k

OpenStudy (jiteshmeghwal9):

volume=\(0.275 m^3\) \(1m^3=10^6cm^3\) \(0.275m^3=0.275 \times 10^6 cm^3\)

OpenStudy (jiteshmeghwal9):

\[1 litre= 1000 cm^3\] \[0.275 \times 10^6 cm^3= ??? litres\]

OpenStudy (cometailcane):

275,000 liters